Your classmate requests you to help him/her in understanding the solutions of the

lim

t→0

(1−cos3t)/

(3t)

= 0, via writing. How do you apply your knowledge based on what you have


learned from this module in doing the task?


1
Expert's answer
2021-05-07T13:29:01-0400

Ans:-

"\lim_{t \to 0} \dfrac{1-Cos3t}{3t}\\
\Rightarrow \lim_{t \to 0} \dfrac{1-(1-2Sin^23t)}{3t}\\
\Rightarrow\lim_{t \to 0} \dfrac{2Sin^23t}{3t}\\
 \\"


"\Rightarrow \lim_{x \to 0} 2\times\dfrac{Sin^23t\times3t}{(3t)^2}\\
\Rightarrow\lim_{t \to 0}6t\lim_{t \to 0} \dfrac{Sin^23t}{(3t)^2}" "\because \lim_{t \to 0} \dfrac{Sint}{t}=1"


"\Rightarrow 
\lim_{t \to 0}6t \times1" "\because \lim_{t \to 0} \dfrac{Sin^23t}{(3t)^2}=1"


"\Rightarrow 0" .....Ans